OpenStudy (anonymous):

A 10−pound package of cookies costs (y + 1) dollars while a 6−pound package costs (y − 1) dollars. If the price per pound of cookies is the same for both packages, the equation below can be used to solve for y. y+1/10 = y-1/6 What is the value, in dollars, of y?

OpenStudy (anonymous):

@Callisto @.Sam. @AccessDenied @dpaInc @myininaya

OpenStudy (callisto):

\[\frac{y+1}{10} = \frac{y-1}{6}\]\[6(y+1)= 10(y-1)\]6y+6 = 10y -10 4y= 16 y=?

OpenStudy (anonymous):

4

OpenStudy (callisto):

Yes !~ :)

OpenStudy (anonymous):

Thanks so much!